OpenStudy (anonymous):

determine 2 coterminal angles one pos one neg for theta =14pi/15 .how is this solved?

OpenStudy (anonymous):

add 2pi for positive then subtract 2pi for negative

OpenStudy (anonymous):

so \[+ \theta = \frac{14\pi}{15} + \frac{30\pi}{15}\] \[-\theta = \frac{14\pi}{15} - \frac{30\pi}{15}\]

OpenStudy (anonymous):

its not suppose to be 14pi/14+2pi/1 = 16pi/15?

OpenStudy (anonymous):

\[\frac{14\pi}{15} + \frac{2\pi}1 \ne \frac{16\pi}{15}\]

OpenStudy (anonymous):

remember your elementary math. you have to convert fractions to common denominators

OpenStudy (anonymous):

\[2\pi = 2\pi \times \frac{15}{15} = \frac{30\pi}{15}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks